2x+4x-2x+5y-6y-8z+15z+z

Answer by Boreal(15235)

2x+4x-2x+5y-6y-8z+15z+z
add like terms
you can add 2x+4x-2x=4x
you can add 5y-6y=-y
you can add -8z+15z+z=8z
the answer is 4x-y+8z